To raise the temperature of 5 kg of water from 20 0C to 30 0C, a 2-kg iron bar is heated and then dropped into the water. What should the temperature of the bar be


1
Expert's answer
2021-04-13T10:54:46-0400

Heat gained = Heat loss

"m _{w}C_{w}\u0394\u03b8_{w} = m _{b}C_{b}\u0394\u03b8_{b}"

where

m = mass ( kg ); mass of water = 5kg, mass of iron bar = 2kg

C = Specific heat capacity; SHC of water = 4182 J/kg°C, SHC of Iron = 462 J/kg°C

Δθ = Temperature change; temperature change in water = 10°C, temperature change in iron bar = "(x- 30)"°C


"5\u00d7 4182\u00d7 10 = 2 \u00d7 462\u00d7 (x- 30)"


"x - 30 = \frac{209100}{924}"


"\therefore x = 256.3 °C"